Labor Force Home Foreign Marginal Product of Labor Shirts Apples MPLs = 2 MPLA = 5 MPLs* = 1 MPLA* = 2 L =100 L* = 240For Questions 9-11 below, assume that the Home and Foreign countries are engaged in free trade. 9. The Foreign country export
Answer #1

Page No - 1 to Answer al Civen that The Intormation provided oppotunity cost of producing (42)/(1/5) = 5/2 too home country.Page No: 2 Here, The PPF of home country is s/2 +AIS = 100! Now, under the tree trade home will com Pletely speciallized ToThe tollowing digs shows the CPF of tre voye No. 3 - home country A Apple C1:200015 - SCO shirt (2= 2500 / - 227.27 Here Cic